OpenStudy (selena2345):

A delivery truck traveled 133 miles in 3.5 hours.What was the average speed of the delivery truck in miles per hour ?

OpenStudy (ashleyisakitty):

Average time=distance/time

OpenStudy (selena2345):

3.5/133=0.026315789

OpenStudy (dmezzullo):

divide 133 into 3.8

OpenStudy (dmezzullo):

Wut do u get then?

OpenStudy (selena2345):

133/3.8=35

OpenStudy (dmezzullo):

I got 38?

OpenStudy (selena2345):

i got 35

OpenStudy (dmezzullo):

U have a typo

OpenStudy (ashleyisakitty):

She put in 3.8, not 3.5. The correct answer is 38.

OpenStudy (dmezzullo):

Is it 3.5 hrs or 3.8 hrs?

OpenStudy (dmezzullo):

Ashley is correct, that is wut i got

OpenStudy (selena2345):

no you told me too divide 133/3.8

OpenStudy (dmezzullo):

Well i am dyslexic

OpenStudy (selena2345):

how did you get 38

OpenStudy (dmezzullo):

so therefore i am thinking one thing and type another

OpenStudy (dmezzullo):

Ok now i aint tryin to be rude or anything but if ur gonna be rude i aint gonna help u again

OpenStudy (selena2345):

Sorry Dmezzullo

OpenStudy (dmezzullo):

Its ok

OpenStudy (selena2345):

you are right @dmezzullo

OpenStudy (dmezzullo):

right about wut?

OpenStudy (selena2345):

133/3.5=38

OpenStudy (dmezzullo):

Okie dokie

OpenStudy (selena2345):

which grade are you in @dmezzullo

OpenStudy (ashleyisakitty):

If you ever divide a number and you get a nonsensical answer like "0.026315789" try and reverse the numbers.

OpenStudy (dmezzullo):

Ya^

OpenStudy (dmezzullo):

I am in 9th

OpenStudy (selena2345):

oh i am in 7th grade

OpenStudy (dmezzullo):

Thats cool. And also, Call me Dmezz or Mezz

OpenStudy (selena2345):

do you live in the U.S @dmezzullo

OpenStudy (dmezzullo):

Yes

OpenStudy (selena2345):

OKAY Dmezz

OpenStudy (selena2345):

Which State Dmezz

OpenStudy (dean.shyy):

Take a look at this: http://is.gd/OkGxHa

OpenStudy (selena2345):

I Live in New York which state do you live in Dmezz

OpenStudy (selena2345):

Dmezzullo

OpenStudy (dmezzullo):

Does tht really Matter?

OpenStudy (ashleyisakitty):

there will be no personal info given out past this point. please close your question, Selena.

OpenStudy (selena2345):

Shut up @Ashleyisakitty

OpenStudy (ashleyisakitty):

Lets let the newbie handle this one, what do you have to say about this @dmezzullo ? :P

OpenStudy (dmezzullo):

@amistre64 @Callisto Take care of this plz

OpenStudy (dmezzullo):

@Selena2345 Do not tell a Ambassador to Shut Up @Selena2345 Ambassadors are here to help users out and keep new and old users outta trouble, So plz be nice and respectful to all users on this site. Thanks and Have a Wonderful Day!

OpenStudy (selena2345):

Sorry @Ashleyisakitty and @dmezzullo

OpenStudy (dmezzullo):

Its ok, But plz in the future don't ask or give out any personal info and Be respectful to any and all users.

OpenStudy (selena2345):

Sorry @dmezzullo